Question:

A particle moves such that its vector is given by π‘Ÿ

= cos πœ”π‘‘π‘– + sin πœ”π‘‘π‘— where πœ” is a constant. Show that:

(i)

Velocity, 𝑣

of the particle is perpendicular to π‘Ÿ


(ii)

Acceleration, π‘Ž

is directed towards the origin and has magnitude proportional to the

distance from the origin.

(iii) π‘Ÿ

𝑋 𝑣

is a constant vector


1
Expert's answer
2021-09-03T07:37:15-0400

The position vector is given by, "\vec{r}=cos\omega t \hat{i}+sin\omega t \hat{j}"


(a) Velocity is given by,

"\vec{v}= \frac{d}{dt}\vec{r}=\frac{d}{dt}(cos\omega t \hat{i}+sin\omega t \hat{j}) = -\omega sin\omega t \hat{i}+\omega cos\omega t \hat{j}"

If velocity is perpendicular to position then, "\vec{r}.\vec{v} = 0"

"\vec{r}.\vec{v} = (cos\omega t \hat{i}+sin\omega t \hat{j}).(-\omega sin\omega t \hat{i}+\omega cos\omega t \hat{j})"

"\vec{r}.\vec{v} = -\omega sin\omega t cos\omega t+\omega sin\omega t cos\omega t = 0"

So, velocity and position are perpendicular.


(b) "\vec{a} = \frac{d}{dt} \vec{ v } =\frac{d}{dt}( -\omega sin\omega t \hat{i}+\omega cos\omega t \hat{j}) = -\omega^2 cos\omega t\hat{i} -\omega^2 sin\omega t\hat{j} = -\omega^2 \vec{r}"

A negative sign indicates that it is acting toward the origin.

"|\vec{a}| = \omega^2|\vec{r}| \implies |\vec{a}| \propto|\vec{r}|"


(c) "\vec{r} \times \vec{v} = (cos\omega t \hat{i}+sin\omega t \hat{j}) \times ( -\omega sin\omega t \hat{i}+\omega cos\omega t \hat{j})"

"\vec{r} \times \vec{v} = (\omega cos^2\omega t+ \omega sin^2\omega t)\hat{k} = \omega \hat{k}"

Hence, it is a constant vector.